Exercise 11-6 (Video) Lewis Company’s standard labor cost of producing one unit of Product DD is 3.40 hours at the rate of $13.80 per hour. During August, 43,700 hours of labor are incurred at a cost of $14.00 per hour to produce 12,600 units of Product DD. (a) Compute the total labor variance. Total labor variance $ (b) Compute the labor price and quantity variances. Labor price variance $ Labor quantity variance $ (c) Compute the labor price and quantity variances, assuming the standard is 3.80 hours of direct labor at $14.20 per hour. Labor price variance $ Labor quantity variance $
Answer #1
(a)
Total labor variance $20,608 Adverse (Unfavorable)
Working:
Total labor variance = Standard labor cost -Actual Labor cost
= (12600*3.4*13.8)-(43700*14)
=          -20,608
(b)
Labor Price Variance $8,740 Adverse (Unfavorable)
Labor quantity variance $11,868 Adverse (Unfavorable)
Working:
Labor price variance = (Standard Price-Actual Price)*Actual labor hours
= (13.8-14.00)*43700
= -8740
Labor quantity Variance = (Standard labor hours-Actual labor hours)*Standard Labor rate
= ((12600*3.4)-43700)*13.80
= -11868
(c)
Labor Price Variance $         8,740 Favorable
Labor quantity variance $       59,356 Favorable
Working:
Labor price variance = (Standard Price-Actual Price)*Actual labor hours
= (14.20-14.00)*43700
= 8740
Labor quantity Variance = (Standard labor hours-Actual labor hours)*Standard Labor rate
= ((12600*3.8)-43700)*14.20
= $       59,356